40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 5 20:37:34 2011, jamie, Summary, General, Summary of today's in-vacuum work 
    Reply  Sat Aug 6 03:10:05 2011, Suresh, Summary, General, Summary of today's in-vacuum work ETMX_OSEMs.pngAfter_osem_adjustment.JPGBefore_osem_adj.JPG
       Reply  Sat Aug 6 13:38:02 2011, rana, Summary, General, Summary of today's in-vacuum work 
          Reply  Sun Aug 7 11:10:54 2011, Suresh, Summary, General, OSEM positioning 
    Reply  Sun Aug 7 14:01:58 2011, kiwamu, Update, Green Locking, X green beam re-aligned 
       Reply  Sun Aug 7 18:41:12 2011, Jenne, Update, Green Locking, X green beam re-aligned 
Message ID: 5132     Entry time: Sun Aug 7 11:10:54 2011     In reply to: 5131
Author: Suresh 
Type: Summary 
Category: General 
Subject: OSEM positioning 

Quote:

This OSEM placement is just the OPPOSITE of what the proper placement is.

Usually, we want to put them in so that the LED beam is vertical. This makes the OSEM immune to the optic's vertical mode.

The orientation with the horizontal LED beam makes the immunity to the side mode better, but may spoil the vertical.

In reality, neither of these assumptions is quite right. The LED beam doesn't come out straight. That's why Osamu and I found that we have to put in some custom orientations.

Also, the magnet gluings relative to the OSEM bracket centers are not perfectly aligned. So...I am saying that the OSEMs have to be oriented empirically to reduce the couplings which we want to reduce.

 

 Kiwamu suggested that since the side resonance is at a lower frequency than the bounce (~17Hz)  we ought to worry about the side more than the bounce.  If this is okay we can reposition the OSEMs to minimise this coupling. 

More over, in the current position, the OSEM s will not sense the side motion!!  So we definitely need to reposition them.  Sorry! I was being a spatz. 

ELOG V3.1.3-